# Generalizable Human Gaussians from Single-View Image

### v4

@@ Abstract
In this work, we tackle the task of learning 3D human Gaussians from a single image, focusing on recovering detailed appearance and geometry including unobserved regions. We introduce a single-view generalizable Human Gaussian Model (HGM), which employs a novel generate-then-refine pipeline with the guidance from human body prior and diffusion prior. Our approach uses a ControlNet to refine rendered back-view images from coarse predicted human Gaussians, then uses the refined image along with the input image to reconstruct refined human Gaussians. To mitigate the potential generation of unrealistic human poses and shapes, we incorporate human priors from the SMPL-X model as a dual branch, propagating image features from the SMPL-X volume to the image Gaussians using sparse convolution and attention mechanisms. Given that the initial SMPL-X estimation might be inaccurate, we gradually refine it with our HGM model. We validate our approach on several publicly available datasets. Our method surpasses previous methods in both novel view synthesis and surface reconstruction. Our approach also exhibits strong generalization for cross-dataset evaluation and in-the-wild images.
@@ 1 Introduction
Automatic 3D human reconstruction from single image is crucial in augmented and virtual reality (AR/VR), game industry, filmmaking, etc. Previous works rely on strong 3D supervision such as the signed distance value or occupancy pifu19 ; pifuhd20 ; gta23 ; icon22 ; econ23 ; sifu24 ; sith24 and focus on surface reconstruction, neglecting novel view synthesis quality, resulting in smoothed and blurred textures. With the development of neural radiance fields nerf , novel view rendering quality has been greatly improved with for human appearance modelling sherf23 ; nhp21 ; mps22 . However, due to the ill-posed nature of single view reconstruction, the back and side views are always blurry and lack details without additional prior. Furthermore, these methods needs large amounts of query points sampled for volume rendering, which hinders practical real-time application in the industries. Some other methods optimize underlying appearance and geometry from scratch by introducing score distillation sampling during the optimizationdreamgaussian24 ; dreamavatar24 . Although effective, these methods still suffer from long time optimization and over-saturation problems. Recent 3D Gaussians generation methods lgm24 ; grm24 combine multi-view diffusion models 0123 ; imagedream23 ; mvdream23 ; instant3d ; dmv3d23 with generalizable multi-view Gaussians prediction models to generate 3D Gaussians with high quality and efficiency. We aim to extend this on human reconstruction. However, directly employ such methods to human reconstruction with complex texture and poses
@@
gives unsatisfying results due to: (1) Inconsistency across multiple views: The multi-view images generated from diffusion model lacks consistency in appearance and pose across different viewpoints. This inconsistency stems from the inherent complexity of human body structure and movement, which leads to low-quality reconstruction results. (2) Quality loss in front view reconstruction: multi-view diffusion process involves down-sampling and changing the original input image. This step results in significant quality degradation when reconstructing the front view image, compromising the fidelity to the original input. 3) Estimating SMPL-X parameters from single view input is ill-posed, directly applying initially estimated SMPL-X can lead to bending legs and wrong elevation issues in previous methodicon22 ; econ23 ; sifu24 ; sith24 .
To address the above-mentioned problems, we introduce a novel Human Gaussians Model (HGM), which supports fast and high quality rendering from single view input, and generalize well to loosing clothes, challenging poses and in-the-wild images as shown in Fig. 1. We do not use multi-view diffusion models due to the multiview inconsistency and resolution degradation problem. Instead, we propose a coarse-to-fine framework, where the diffusion model is adapted to refine back-view images rendered from our predicted coarse human Gaussians. In this way, we can keep the resolution and content of the original input image for high-fidelity reconstruction. In order to model the complex structure of a human, we inject the human prior into the Gaussian prediction process. Specifically, our model consists of two branches: 1) The first branch is a UNet to directly predict Gaussians from the input image, as inspired by image splatter splatter24 . 2) The second branch uses learnable tokens attached to SMPL-X vertices for structural feature extraction with attention layers, and then combined with UNet features with SparseConvspconv18 and a transformer for Gaussian enhancement. Recognizing the inaccurate estimate of SMPL-X from the pre-trained model, during inference, we iteratively refine the initial SMPL-X parameters with our HGM pre-trained with ground truth SMPL-X. Given the loss of details of the back view by directly predicting the Gaussians from a single view, we further apply a ControlNet to refine the back view with the control signal from the back-view image rendered from the coarse stage. We then input the original front view and refined back view images to our HGM model to get the final refined Gaussians. Meshes can be extracted from densely rendered depth map and TSDF fusion. Our model can be trained with only posed multiview images without 3D supervision and generalizes well to untrained datasets and in the wild images.
@@ 2 Related works
Single-view Human Reconstruction. PIFu pifu19 , PIFuHD pifuhd20 , PaMIR pamir20 , and GTA gta23 are capable of inferring full textures from a single image. Techniques such as PHORHUM phorhum22 and S3F s3f23 go further by separating albedo and global illumination. However, these methods lack information from other views or prior knowledge, such as diffusion models, often resulting in unsatisfactory textures. TeCH tech24 utilizes diffusion-based models to visualize unseen areas, producing realistic results. However, it requires time-intensive optimization per subject and is dependent on accurate SMPL-X. The emergence of Neural Radiance Fields (NeRF) has led to methods sherf23 ; elicit23 ; mps22 ; nhp21 using videos or multi-view images to optimize NeRF for the capture of human forms. Recent advancements like SHERF sherf23 and ELICIT elicit23 aim to generate human NeRFs from single images. Although NeRF-based approaches are effective in creating high-quality images from various perspectives, they often struggle with detailed 3D mesh generation from single images and require extensive optimization time. More recently, SiTH sith24 proposes to combine a back-view hallucination model with an SDF-based mesh reconstruction model. Similarly, SIFU sifu24 employs a text-to-image diffusion-based prior to generating consistent textures for invisible views. However, these methods require 3D annotations such as the SDF of the meshes and texture maps as strong supervision and still fail to generate renderings with high fidelity due to the limited 3D training data and representation. In addition, these methods suffer from SMPL estimation errors, leading to bending legs and wrong elevation of the reconstructed 3D humans. Compared to these methods, our approach can be trained solely on multi-view images and achieves much better novel view synthesis quality.
Human Gaussians. 3D Gaussians gs23 and differentiable splatting splatter24 have gained broad popularity due to their efficiency in reconstructing high-fidelity 3D scenes from posed images using only a moderate number of 3D Gaussians. This representation has been quickly adopted for various applications, including imag or text-conditioned 3D generation and avatar reconstruction. Among these methods, Gauhuamn and HUGS gauhuman ; hugs are the first to propose optimizing human Guassians from monocular human videos. However, they are not applicable to single static human images. GPS-Gaussiangpsgaussian propose a generalizable multi-view huaman Gaussian model with high quality rendering; however, it needs dense views 16 or 8, which cannot be directly applied to single-view human images. In contrast, our human model is able to generate human Gaussians from single view image with strong generalization.
@@
Generalizable Gaussians with Multi-view Diffusion. The Large Reconstruction Model (LRM) lrm24 scales up both the model and the dataset to predict a neural radiance field (NeRF) from single-view images. Although LRM is primarily a reconstruction model, it can be combined with Diffusion Models (DMs) to achieve text-to-3D and image-to-3D generation as demonstrated by extensions such as Zero1230123 , Image Dreamimagedream23 Instant3D instant3d and DMV3D dmv3d23 . Our method also builds on a strong reconstruction model and uses pre-trained 2D DMs to provide input images missing information in a feedforward manner. Some concurrent works, such as LGM lgm24 , AGG agg24 , and Splatter Image splatter24 , also utilize 3D Gaussians in a feed-forward model. LGM lgm24 combines novel view generation diffusion models with generalizable Gaussians in a feedforward manner, while GRM grm24 replaces the U-Net architecture with a pure-transformer one and scales up to large resolution. However, these methods face two main challenges when using pre-trained diffusion models. Firstly, the generated input view image becomes blurry compared to the original input, which affects the subsequent generalizable Gaussian model. Secondly, diffusion models can introduce multiview inconsistency, especially for human images with different poses, making direct adaptation unfeasible. We solve these problems by using ControlNet as the refinement tools without damaging the input image quality or introducing multi-view inconsistency.
@@ 3.1 Preliminaries
3D Gaussian Splatting (3DGS). Introduced by gs23 , 3D Gaussian splatting represents 3D assets or scenes using a collection of 3D Gaussians. Each Gaussian is characterized by its center $x\in\mathbb{R}^{3}$, scaling factor $s\in\mathbb{R}^{3}$, rotation $r\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}$, and color features $c\in\mathbb{R}^{c}$. View-dependent effects can be modeled with spherical harmonics. 3D scenes can be explicitly represented by a set of Gaussians $G=\{G_{i}\}$, where $G_{i}=\{{x_{i},s_{i},r_{i},\alpha_{i},c_{i}}\}$ represents the attributes for the $i$-th Gaussian. Compared with NeRF nerf , 3DGS performs fast rendering by first projecting Gaussians onto the image plane as 2D Gaussians and performing alpha-blending for each pixel in front-to-back depth order. Building on this, Image Splatter splatter24 proposes predicting Gaussians from a single image through image-to-image translation. Specifically, each pixel is converted to a Gaussian with corresponding attributes, supervised by multi-view images. Our model builds on this representation by directly predicting XYZ coordinates from the image instead of the depth.
@@ 3.3.1 Our HGM model
The lower part of Fig. 2 shows our proposed Human Gaussian Model (HGM). The direct prediction of Gaussians from image pixels with UNet splatter24 ; lgm24 lacks human shape and pose prior, thus leading to unsatisfactory results. We therefore introduce a dual branch that utilizes SMPL-X to enforce human shape and pose prior into the Gaussian prediction process. Specifically, for UNet branch, the collection of the RGB value and ray embedding for each pixel are concatenated into a 9-channel feature map as the input $F_{I}=\{c_{i},o_{i}\times d_{i},d_{i}\mid i=1,2,...,N\}$. Our HGM model predicts Gaussians from the U-Net as:
@@
We call it as Image Gaussians. For the SMPL-X branch, we attach learnable tokens to each of the SMPL-X vertices and extract the patch features of the image, denoted $F^{\prime}_{I}$. We use cross-attention between these learnable tokens and the patch features to obtain $F_{S}$. This approach takes advantage of the fact that SMPL-X vertices are defined in semantically similar areas across different identities. Consequently, the learned tokens can memorize the mapping from the training dataset to unseen identities during inference, effectively providing structural human priors. This mechanism enables our model to capture and utilize semantic consistent features across diverse identities, enhancing its ability to generalize to new subjects. We apply SparseConvNet spconv18 (SPConv $\Psi$) to propagate the SMPL-X features to the predefined whole bounding box, and we denote this feature as the SMPL-X volume feature: $S^{vol}=\Psi(F_{S})$, where $F_{S}$ are the SMPL-X features. The volume feature reconstructed from the SMPL-X vertex feature provides geometric cues of the target human body. The centers of the Image Gaussians from $G_{u}$ are then used to sample the propagated SMPL-X volume features, denoted as $f_{s}=S^{vol}(C_{u})$, where $C_{u}$ are the centers of $G_{u}$. SMPL-X $f_{s}$ features are then concatenated with the features of UNet $f_{u}$ for each Gaussian. This concatenated feature is fed into a transformer $\operatorname{Tr}_{mix}$ to obtain the coarse Gaussians shown in 2. Specifically, we predict the xyz coordinates residuals for the Image Gaussians and all the other updated Gaussian features. $\operatorname{Tr}_{mix}$ is a transformer that contains multiple self-attention blocks among Gaussians to ensure that each Gaussian is aware of the other.
@@ 3.3.2 SMPL-X Refinement
Given initial estimated SMPL-X is not accurate, we leverage our pre-trained HGM to iteratively refine the SMPL-X parameters based on the mask and normal matching as shown in Fig. 3 left part. Specifically, we use the initial estimated SMPL-X from pre-trained SMPL-X estimator to reconstruct the initial Gaussians and rendered the side-view masks and normals. We minimize the mask and normal difference between the Gaussian and SMPL-X renderings and back-propagate the loss to SMPL-X parameters. Then we interatively input the updated SMPL-X to our HGM model, so the Gaussian is also updated to give more accurate masks. For normal matching, we use pre-trained normal estimator from icon22 which also needs SMPL-X as input and iteratively update the SMPL-X parameters. Specifically, we compute the $\mathcal{L}_{SMPL-X}$ as shown in 3 and back propagate it to SMPL-X parameters. $\mathcal{L}_{normal}$, $\mathcal{L}_{front}$ and $\mathcal{L}_{side}$ are $\mathcal{L}_{1}$ losses between the rendered masks of SMPL-X model and coarse Gaussians by HGM. $\mathcal{L}_{normal}$ is the $\mathcal{L}_{1}$ loss between the rendered SMPLX normal and the pre-trained normal estimator icon22 . Note that both of the normal and side mask supervision are updated as the input of the HGM and normal estimator also contains the updated SMPL-X in each iteration. The whole process takes 100 iterations. We provide more analysis on our SMPL-X refinement in the appendix.
$\mathcal{L}_{SMPL-X}=\lambda_{front}\mathcal{L}_{m_{front}}+\lambda_{side}\mathcal{L}_{m_{side}}+\lambda_{n}\mathcal{L}_{normal}$ (3)
Figure 4: Our back-view refinement can generate more realistic back-view images, compared with back-view hallucination of SiTH sith24 .

### v5

@@ Abstract
In this work, we tackle the task of learning 3D human Gaussians from a single image, focusing on recovering detailed appearance and geometry including unobserved regions. We introduce a single-view generalizable Human Gaussian Model (HGM), which employs a novel generate-then-refine pipeline with the guidance from human body prior and diffusion prior. Our approach uses a ControlNet to refine rendered back-view images from coarse predicted human Gaussians, then uses the refined image along with the input image to reconstruct refined human Gaussians. To mitigate the potential generation of unrealistic human poses and shapes, we incorporate human priors from the SMPL-X model as a dual branch, propagating image features from the SMPL-X volume to the image Gaussians using sparse convolution and attention mechanisms. Given that the initial SMPL-X estimation might be inaccurate, we gradually refine it with our HGM model. We validate our approach on several publicly available datasets. Our method surpasses previous methods in both novel view synthesis and surface reconstruction. Our approach also exhibits strong generalization for cross-dataset evaluation and in-the-wild images. We open-source our code at: https://github.com/jinnan-chen/HGM.
@@ 1 Introduction
Automatic 3D human reconstruction from single image is crucial in augmented and virtual reality (AR/VR), game industry, filmmaking, etc. Previous works rely on strong 3D supervision such as the signed distance value or occupancy (Saito et al., 2019; 2020; Zhang et al., 2023c; Xiu et al., 2022; 2023; Zhang et al., 2024; Ho et al., 2024) and focus on surface reconstruction, neglecting novel view synthesis quality, resulting in smoothed and blurred textures. With the development of neural radiance fields (Mildenhall et al., 2020), novel view rendering quality has been greatly improved for human appearance modeling (Hu et al., 2023; Kwon et al., 2021; Gao et al., 2022). However, due to the ill-posed nature of single view reconstruction, the back and side views are always blurry and lack details without additional prior. Furthermore, these methods needs large amounts of query points sampled for volume rendering, which hinders practical real-time application in the industries. Some other methods optimize underlying appearance and geometry from scratch by introducing score distillation sampling during the optimizationTang et al. (2024b); Cao et al. (2024). Although effective, these methods still suffer from slow optimization and over-saturation problems. Recent 3D Gaussians generation methods (Tang et al., 2024a; Yinghao et al., 2024) combine multi-view diffusion models (Liu et al., 2023; Wang & Shi, 2023; Shi et al., 2023; Li et al., 2024; Xu et al., 2023) with generalizable multi-view Gaussians prediction models to generate 3D Gaussians with high quality and efficiency. We aim to extend this on human reconstruction. However, directly employ such methods to human reconstruction with complex texture and poses
@@
gives unsatisfying results due to: 1) Inconsistency across multiple views: The multi-view images generated from diffusion model lacks consistency in appearance and pose across different viewpoints. This inconsistency stems from the inherent complexity of human body structure and movement, which leads to low-quality reconstruction results. 2) Quality loss in front view reconstruction: multi-view diffusion process involves down-sampling and changing the original input image. This step results in significant quality degradation when reconstructing the front view image, compromising the fidelity to the original input. 3) Estimating SMPL-X parameters from single view input is ill-posed, directly applying initially estimated SMPL-X can lead to bending legs and wrong elevation issues in previous method Xiu et al. (2022; 2023); Zhang et al. (2024); Ho et al. (2024).
To address the above-mentioned problems, we introduce a novel Human Gaussians Model (HGM), which supports fast and high quality rendering from single view input, and generalize well to loosing clothes, challenging poses and in-the-wild images as shown in Fig. 1. We do not use multi-view diffusion models due to the multiview inconsistency and resolution degradation problem. Instead, we propose a coarse-to-fine framework, where the diffusion model is adapted to refine back-view images rendered from our predicted coarse human Gaussians. In this way, we can keep the resolution and content of the original input image for high-fidelity reconstruction. In order to model the complex structure of a human, we inject the human prior into the Gaussian prediction process. Specifically, our model consists of two branches: 1) The first branch is a UNet to directly predict Gaussians from the input image, as inspired by image splatter (Szymanowicz et al., 2024). 2) The second branch uses learnable tokens attached to SMPL-X vertices for structural feature extraction with attention layers, and then combined with UNet features with SparseConv Graham et al. (2018) and a transformer for Gaussian enhancement. Recognizing the inaccurate estimate of SMPL-X from the pre-trained model, during inference, we iteratively refine the initial SMPL-X parameters with our HGM pre-trained with ground truth SMPL-X. Given the loss of details of the back view by directly predicting the Gaussians from a single view, we further apply a ControlNet to refine the back view with the control signal from the back-view image rendered from the coarse stage. We then input the original front view and refined back view images to our HGM model to get the final refined Gaussians. Meshes can be extracted from densely rendered depth map and TSDF fusion. Our model can be trained with only posed multiview images without 3D supervision and generalizes well to untrained datasets and in-the-wild images.
@@ 2 Related works
Single-view Human Reconstruction. PIFu (Saito et al., 2019), PIFuHD (Saito et al., 2020), PaMIR (Zheng et al., 2021), and GTA (Zhang et al., 2023c) are capable of inferring full textures from a single image. Techniques such as PHORHUM (Alldieck et al., 2022) and S3F (Corona et al., 2023) go further by separating albedo and global illumination. However, these methods lack information from other views or prior knowledge, such as diffusion models, often resulting in unsatisfactory textures. TeCH (Huang et al., 2024) utilizes diffusion-based models to visualize unseen areas, producing realistic results. However, it requires time-intensive optimization per subject and is dependent on accurate SMPL-X. The emergence of Neural Radiance Fields (NeRF) has led to methods (Hu et al., 2023; Huang et al., 2023; Gao et al., 2022; Kwon et al., 2021) using videos or multi-view images to optimize NeRF for the capture of human forms. Recent advances such as SHERF (Hu et al., 2023) and ELICIT (Huang et al., 2023) aim to generate human NeRFs from single images. Although NeRF-based approaches are effective in creating high-quality images from various perspectives, they often struggle with detailed 3D mesh generation from single images and require extensive optimization time. More recently, SiTH (Ho et al., 2024) proposes to combine a back-view hallucination model with an SDF-based mesh reconstruction model. Similarly, SIFU (Zhang et al., 2024) employs a text-to-image diffusion-based prior to generating consistent textures for invisible views. However, these methods require 3D annotations such as the SDF of the meshes and texture maps as strong supervision and still fail to generate renderings with high fidelity due to the limited 3D training data and representation. In addition, these methods suffer from SMPL estimation errors, leading to bending legs and wrong elevation of the reconstructed 3D humans. Compared to these methods, our approach can be trained solely on multi-view images and achieves much better novel view synthesis quality. Human Gaussians. 3D Gaussians (Kerbl et al., 2023) and differentiable splatting (Szymanowicz et al., 2024) have gained broad popularity due to their efficiency in reconstructing high-fidelity 3D scenes from posed images using only a moderate number of 3D Gaussians. This representation has been quickly adopted for various applications, including imag or text-conditioned 3D generation and avatar reconstruction. Among these methods, Gauhuamn and HUGS (Hu & Liu, 2024; Kocabas et al., 2024) are the first to propose optimizing human Guassians from monocular human videos. However, they are not applicable to single static human images. GPS-Gaussian(Zheng et al., 2024) propose a generalizable multi-view huaman Gaussian model with high quality rendering; however, it needs dense views 16 or 8, which cannot be directly applied to single-view human images. Our human model achieves strong generalization in generating human Gaussians from single-view images, complementing concurrent work such as Pan et al. (2024).
@@
Generalizable Gaussians with Multi-view Diffusion. The Large Reconstruction Model (LRM) (Hong et al., 2024) scales up both the model and the dataset to predict a neural radiance field (NeRF) from single-view images. Although LRM is primarily a reconstruction model, it can be combined with Diffusion Models (DMs) to achieve text-to-3D and image-to-3D generation as demonstrated by extensions such as Zero123(Liu et al., 2023), Image Dream(Wang & Shi, 2023) Instant3D (Li et al., 2024) and DMV3D (Xu et al., 2023). Our method also builds on a strong reconstruction model and uses pre-trained 2D DMs to provide input images missing information in a feedforward manner. Some concurrent works, such as LGM (Tang et al., 2024a), AGG (Xu et al., 2024), and Splatter Image (Szymanowicz et al., 2024), also utilize 3D Gaussians in a feed-forward model. LGM (Tang et al., 2024a) combines novel view generation diffusion models with generalizable Gaussians in a feedforward manner, while GRM (Yinghao et al., 2024) replaces the U-Net architecture with a pure-transformer one and scales up to large resolution. However, these methods face two main challenges when using pre-trained diffusion models. Firstly, the generated input view image becomes blurry compared to the original input, which affects the subsequent generalizable Gaussian model. Secondly, diffusion models can introduce multiview inconsistency, especially for human images with different poses, making direct adaptation unfeasible. We solve these problems by using ControlNet as the refinement tools without damaging the input image quality or introducing multi-view inconsistency.
@@ 3.1 Preliminaries
3D Gaussian Splatting (3DGS). Introduced by (Kerbl et al., 2023), 3D Gaussian splatting represents 3D assets or scenes using a collection of 3D Gaussians. Each Gaussian is characterized by its center $x\in\mathbb{R}^{3}$, scaling factor $s\in\mathbb{R}^{3}$, rotation $r\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}$, and color features $c\in\mathbb{R}^{c}$. View-dependent effects can be modeled with spherical harmonics. 3D scenes can be explicitly represented by a set of Gaussians $G=\{G_{i}\}$, where $G_{i}=\{{x_{i},s_{i},r_{i},\alpha_{i},c_{i}}\}$ represents the attributes for the $i$-th Gaussian. Compared with NeRF (Mildenhall et al., 2020), 3DGS performs fast rendering by first projecting Gaussians onto the image plane as 2D Gaussians and performing alpha-blending for each pixel in front-to-back depth order. Building on this, Image Splatter (Szymanowicz et al., 2024) proposes predicting Gaussians from a single image through image-to-image translation. Specifically, each pixel is converted to a Gaussian with corresponding attributes, supervised by multi-view images. Our model builds on this representation by directly predicting XYZ coordinates from the image instead of the depth.
@@ 3.3.1 Our HGM model
The lower part of Fig. 2 shows our proposed Human Gaussian Model (HGM). The direct prediction of Gaussians from the image pixels with UNet (Szymanowicz et al., 2024; Tang et al., 2024a) lacks human shape and pose prior, thus leading to unsatisfactory results. We therefore introduce a dual branch that utilizes SMPL-X to enforce human shape and pose prior to the Gaussian prediction process. Specifically, for UNet branch, the collection of the RGB value and ray embedding for each pixel are concatenated into a 9-channel feature map as the input $F_{I}=\{c_{i},o_{i}\times d_{i},d_{i}\mid i=1,2,...,N\}$. Our HGM model predicts Gaussians from the U-Net as:
@@
which we refer to as Image Gaussians. For the SMPL-X branch, we attach learnable tokens to each of the SMPL-X vertices and extract the patch features of the image, denoted $F^{\prime}_{I}$. We use cross-attention between these learnable tokens and the patch features to obtain $F_{S}$. This approach takes advantage of the fact that SMPL-X vertices are defined in semantically similar areas across different identities. Consequently, the learned tokens can memorize the mapping from the training dataset to unseen identities during inference, effectively providing structural human priors. This mechanism enables our model to capture and utilize semantic consistent features across diverse identities, enhancing its ability to generalize to new subjects. We apply SparseConvNet (Graham et al., 2018) (SPConv $\Psi$) to propagate the SMPL-X features to the predefined whole bounding box, and we denote this feature as the SMPL-X volume feature: $S^{vol}=\Psi(F_{S})$, where $F_{S}$ are the SMPL-X features. The volume feature reconstructed from the SMPL-X vertex feature provides geometric cues of the target human body. The centers of the Image Gaussians from $G_{u}$ are then used to sample the propagated SMPL-X volume features, denoted as $f_{s}=S^{vol}(C_{u})$, where $C_{u}$ are the centers of $G_{u}$. SMPL-X $f_{s}$ features are then concatenated with the features of UNet $f_{u}$ for each Gaussian. This concatenated feature is fed into a transformer $\operatorname{Tr}_{mix}$ to obtain the coarse Gaussians, i.e.:
@@
Specifically, we predict the xyz coordinates residuals for the Image Gaussians and all the other updated Gaussian features. $\operatorname{Tr}_{mix}$ is a transformer that contains multiple self-attention blocks among Gaussians to ensure that each Gaussian is aware of the other.
@@ 3.3.2 SMPL-X Refinement
Given initial estimated SMPL-X is not accurate, we leverage our pre-trained HGM to iteratively refine the SMPL-X parameters based on the mask and optionally normal matching as shown in Fig. 3 left part. Specifically, we rendered the side-view masks and normals(optionally). We minimize the mask and normal difference between the Gaussian and SMPL-X renderings and back-propagate the loss to SMPL-X parameters. Then we interatively input the updated SMPL-X to our HGM model, so the Gaussian is also updated to give more accurate masks. For normal matching, we use the pre-trained normal estimator from Xiu et al. (2022) that also needs SMPL-X as input and iteratively update the SMPL-X parameters. Specifically, we compute:
$\mathcal{L}_{SMPL-X}=\lambda_{front}\mathcal{L}_{{front}}+\lambda_{side}\mathcal{L}_{{side}}+\lambda_{n}\mathcal{L}_{normal},$ (3)
and back propagate it to the SMPL-X parameters. $\mathcal{L}_{normal}$, $\mathcal{L}_{front}$ and $\mathcal{L}_{side}$ are losses between the rendered masks and 2D detected keypoints of SMPL-X model and the coarse Gaussians by HGM. $\mathcal{L}_{normal}$ is the loss between the rendered SMPLX normal and the pre-trained normal estimator (Xiu et al., 2022). Note that both the normal and side mask supervisions are updated as the input of the HGM and the normal estimator also contains the updated SMPL-X in each iteration. We provide more analysis on our SMPL-X refinement in the appendix.
Figure 4: Our back-view refinement can generate more realistic back-view images, compared with back-view hallucination of SiTH Ho et al. (2024).